OpenStudy (anonymous):

Can anyone help with synthetic division please........ (3x^3+6x^2+8x+6x)/(x-1/2)............. I don't understand synthetic division, or longhand division please please please help me

OpenStudy (anonymous):

Check the ast term. Is it 6 or 6x?

OpenStudy (anonymous):

its 6

OpenStudy (anonymous):

I've been working this problem for 30 min now

OpenStudy (anonymous):

ok. First arrange the coefficients of the given polynomial in descending order Note: for example: 5x^2 - 1 will have the coefficient 5 0 -1

OpenStudy (anonymous):

Next solve the given divisor for its zero

OpenStudy (anonymous):

this is the answer that I get for the problem 3x^2+7x+15/2+13/4(x-1/2)

OpenStudy (anonymous):

So (3x^3 + 6x^2 + 8x + 6) /(x -1/2) will become the following: x - 1/2 = 0 will give x = 1/2 Coefficient for 3x^3 + 6x^2 + 8x + 6 will be 3 6 8 6 Synthetic division will now begin as follows: 1/2 | 3 6 8 6 | 3/2 15/4 47/8 3 15/2 47/4 95/8 Thus, (3x^3 + 6x^2 + 8x + 6) = (3x^2 + (15/2)x + (47/4))(x -1/2) + (95/8)

OpenStudy (anonymous):

it's have to be written as a fraction at the end

OpenStudy (anonymous):

I am typing it out well for you. Will do it as an attachment.

OpenStudy (anonymous):

this is what I get when I check the answer......... Remember to place the remainder over the divisor, them simplify so that there is no fraction in the numerator

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you check on it now?